OpenStudy (anonymous):

Solve: 2(3x+2)2=5(3x+2)-3

OpenStudy (anonymous):

Are you missing something? Because \[2(3x+2)2 = 4(3x+2)\] Did you mean \[2(3x+2) ± 2\]

OpenStudy (anonymous):

anyway… assuming that is correct, we can multiply through: \[2(3x+2)2 = 4(3x+2) = 12x + 8\] and \[5(3x+2)-3 = 15x + 10 - 3 = 15x + 7\] so \[12x + 8 = 15x + 7\] Subtract 12x from both sides, and 7 from both sides, and we get: \[1 = 3x\] So just divide both sides by 3 to get x by itself: \[x = 1/3\]